# The Surprising Effectiveness of Negative Reinforcement in LLM Reasoning

### v1

@@ 4.3 Extending Gradient Analysis to PPO and GRPO
where $A$ denotes the advantage. Comparing the PPO and GRPO objectives to Equation 6, there are three key differences: (1) policy loss clipping, (2) KL regularization (reward penalty in PPO, loss in GRPO), and (3) advantages instead of raw rewards. We analyze their impact on the gradient dynamics below:
@@
Therefore, while PPO and GRPO introduce modifications that stabilize learning, the core gradient behaviors identified in our previous analysis still hold. We further discuss the differences between RLVR and RL with reward models in Appendix B.
@@ 5 Balancing Positive and Negative Reinforcement
We apply $\lambda=0.1$ in our experiments. As shown in Table 1, W-REINFORCE consistently delivers a favorable trade-off across a range of $k$ values on MATH, AIME 2025 and AMC23. On MATH, W-REINFORCE matches the best Pass@$1$ score $76.6$ with PPO and has the highest performance for all $k\leq 64$, while still preserves competitive performance at $k=256$. On AIME 2025, W-REINFORCE performs even more strongly—achieving the best result across all $k$ values except $k=1$. These results confirm that W-REINFORCE, a simple weighted extension of the classic REINFORCE algorithm, strikes an effective balance between the strengths of PSR and NSR. By merely scaling down the weight of positive rewards, it achieves strong performance while preserving diversity. Despite its simplicity, W-REINFORCE consistently outperforms strong RL algorithms such as PPO and GRPO across most Pass@$k$. These findings suggest that this simple variant of REINFORCE can serve as a competitive alternative to more complex RL algorithms when the model prior is strong (e.g., Qwen models).
@@ 6 Related Work
Reinforcement learning has shown great promise in improving large language models, as demonstrated by the success of reinforcement learning from human feedback (RLHF) and from AI feedback (RLAIF), which align model responses with human preferences (lee2024rlaif, ; meng2024simpo, ; ouyang2022training, ; rafailov2023direct, ). More recently, reinforcement learning with verifiable rewards (dai2025s, ; gao2024designing, ; havrilla2024teaching, ; lambert2024t, ; liu2025understanding, ; shao2024deepseekmath, ; team2025intellect, ; wang2025reinforcement, ; xu2025scalable, ; yu2025dapoopensourcellmreinforcement, ; yuan2025vapo, ; yuan2025s, ; zhang2025srpo, ) has attracted growing attention for its effectiveness in incentivizing reasoning in LLMs with rule-based rewards (fatemi2025concise, ; guo2025deepseek, ; kazemnejad2024vineppo, ; team2025kimi, ; zeng2025simplerl, ; dong2024rlhf, ; lambert2024rewardbench, ; mroueh@understandgrpo, ). Notably, DeepSeek-R1 (guo2025deepseek, ) and Kimi K1.5 (team2025kimi, ) demonstrate that RLVR can elicit emergent reasoning behaviors such as long chain of thought and self-reflection, and achieve strong performance across diverse reasoning tasks such as math and coding problems. Yeo et al. (yeo2025demystifying, ) further explore the emergence of long CoT across different RLVR setups.
Despite these advances, many prior works mainly focus on evaluating the model’s Pass@$1$ or greedy decoding performance, which might overlook the underlying change in model behavior (e.g., inference scaling performance). More importantly, the mechanisms behind RLVR for driving reasoning and generalization remain underexplored. In this work, we take a step further by decomposing the RLVR objective into two distinct learning paradigms—learning from correct responses and learning from incorrect responses—and analyzing the learning signal at the token level through gradient analysis. We evaluate the models extensively using Pass@$k$ with a wide spectrum of $k$. Our findings highlight the critical yet previously underappreciated role of negative rewards in RLVR.
@@
Inference-time scaling has emerged as a promising direction for enhancing model performance (balachandran2025inference, ; brown2024large, ; chen2024more, ; lightman2024lets, ; muennighoff2025s1, ; qu2025optimizing, ; setlur2025scaling, ; snell2024scaling, ; wang2024openr, ; wang2024math, ; wang2023selfconsistency, ; welleck2024decoding, ; wu2025inference, ; core23zhu, ; yao2023tree, ; zelikman2022star, ; zhao2025sample, )—particularly in reasoning tasks where generating multiple candidate solutions (cobbe2021training, ; core23zhu, ) or longer reasoning traces (muennighoff2025s1, ; yeo2025demystifying, ; zeng2025revisiting, ) can help hit the correct answer. Through the lens of inference-time scaling, recent studies have raised questions about whether RL-trained models are better than the base models (yue2025does, ; hochlehnert2025sober, ). A recent work by (yue2025does, ) suggests that RLVR primarily adjusts the model’s output distribution toward high-reward responses, rather than eliciting new reasoning abilities beyond the base model. By adopting Pass@$k$ as metric, they find that RL-trained models have inferior inference-time scaling performance than the base models. These findings align with our findings that reinforcing correct samples hurts Pass@$k$ at larger $k$. Another concurrent work by (dang2025weight, ) studies supervised fine-tuning (SFT), and finds that diversity collapse during SFT adversely affects inference-time scaling performance. They show that SFT reduces generation diversity, leading to degraded Pass@$k$ performance, and that interpolating weights from early, potentially less overconfident checkpoints and later ones can effectively restore diversity and improve both Pass@$1$ and Pass@$k$.
@@ 7 Conclusion
In this work, we investigate the mechanism underlying RLVR for LM reasoning. By decomposing RLVR into positive and negative sample reinforcement, we reveal a surprising finding: solely penalizing incorrect samples can effectively enhance LM reasoning capabilities while preserving generation diversity. Experimental results show that NSR consistently improves performance across a wide Pass@$k$ spectrum and in many cases matches or outperforms strong RL algorithms such as PPO and GRPO. Our gradient analysis demonstrates that NSR works by suppressing incorrect responses and redistributing probability mass toward plausible alternatives based on the model prior. Building on these findings, we proposed a simple variant of REINFORCE, Weighted-REINFORCE, that upweights the negative sample reinforcement. Empirical results show that it achieves a good balance between PSR and NSR, and yields consistent Pass@$k$ improvements across multiple reasoning benchmarks. We discuss limitations and future work directions in Appendix D.
@@ Acknowledgments
The authors would like to thank the members of the Princeton NLP Group for their valuable feedback and discussions. This research is partially funded by the National Science Foundation (IIS-2211779), the NVIDIA Academic Grant, the Lambda Research Grant, and the OpenAI Superalignment Fast Grant. MX is supported by the Apple Scholars in AI/ML PhD Fellowship.
@@ Appendix A Gradient Derivation
In summary, the gradient of Equation 6 can be expressed as
$\frac{\partial\mathcal{L}}{\partial z_{v}}=\begin{cases}-R\cdot\pi_{v}\cdot(1-\pi_{v})&\text{if }v=y_{t}\quad\text{(sampled token)}\\ R\cdot\pi_{v}\cdot\pi_{y_{t}}&\text{if }v\neq y_{t}\quad\text{(unsampled token)}\end{cases}$
@@
Appendix B RLVR vs. RL with Reward Models
Different from RLVR, many RL setups require training reward models to assign scalar values that indicate the quality of generation (e.g., in RLHF [32]). During policy optimization, these rewards are often normalized within a batch by subtracting the mean, so that a sample’s final reward depends on its relative ranking compared to others in the same batch. As a result, a sample’s reward can change sign (from positive to negative or vice versa) depending on the overall quality of the batch, making it difficult to interpret rewards as absolute indicators of positive or negative samples. In contrast, RLVR applies a binary reward ($+1$/$-1$). The signs of the rewards are inherently tied to the correctness of individual sequences, as each token in a sequence receives the same reward and the batch average reward always remains within $[-1,1]$, thus reward normalization will preserve the original sign. This inherent sign preservation in RLVR is critical for our analysis, as it enables a clean and unambiguous separation into positive and negative learning paradigms based on the reward signs.
Appendix C Implementation Details
C.1 Objectives and Training Hyperparameters
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\beta\text{KL}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg{)},$
$\mathcal{L}_{\text{PSR}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\min\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},\text{clip}\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},1-\epsilon,1+\epsilon\right)\right),\quad r(\boldsymbol{x},\boldsymbol{y})=1,$
$\mathcal{L}_{\text{NSR}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\min\left(-\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},-\text{clip}\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},1-\epsilon,1+\epsilon\right)\right),\quad r(\boldsymbol{x},\boldsymbol{y})=-1,$
@@
C.2 Prompt Templates
We adopt the prompt templates for Qwen models following [66], as shown in Tables 2 and 3.
@@ C.2 Prompt Templates
C.3 Advantage Normalization
The implementation of computing advantage in verl includes an advantage normalization. However, for PSR and NSR, since we exclude the KL penalty, the advantage is equal to the raw reward (i.e., $+1$ for PSR and $-1$ for NSR). In this case, applying normalization would cause the advantage to be zero, thus eliminating the learning signal. As a result, we disable this normalization in PSR and NSR’s implementation. A recent work [53] also suggests that the reward normalization may not be necessary for certain settings.
Appendix D Limitations
@@
    Primary focus on Qwen models. To understand how PSR and NSR interact with and leverage strong model priors, our experiments concentrate on Qwen2.5 and Qwen3 models that are known for their superior reasoning capabilities. As different LMs possess varying levels of prior knowledge and inductive biases, an interesting future direction is to investigate the learning dynamics of PSR, NSR, and W-REINFORCE across a broader range of model families.
@@
    Performance degradation with extended NSR training. We observe that extensive training with NSR (e.g., over hundreds of gradient steps) leads to a noticeable decline in performance, whereas W-REINFORCE does not exhibit this issue. This suggests that NSR’s implicit mechanism for preserving prior knowledge may be insufficient to ensure stable training over the long term, and incorporating some degree of PSR may be necessary. Future work could explore more sophisticated methods for integrating NSR and PSR to design more robust RL algorithms, and investigate NSR as a potential warm-up phase before standard RL training given that it improves the full Pass@$k$ spectrum.
  3. 3.

### v2

@@ 4.3 Extending Gradient Analysis to PPO and GRPO
where $A$ denotes the advantage. Comparing the PPO and GRPO objectives to Equation˜6, there are three key differences: (1) policy loss clipping, (2) KL regularization (reward penalty in PPO, loss in GRPO), and (3) advantages instead of raw rewards. We analyze their impact on the gradient dynamics below:
@@
Therefore, while PPO and GRPO introduce modifications that stabilize learning, the core gradient behaviors identified in our previous analysis still hold. We further discuss the differences between RLVR and RL with reward models in Appendix˜C.
@@ 5 Balancing Positive and Negative Reinforcement
We apply $\lambda=0.1$ in our experiments. As shown in Table 1, W-REINFORCE consistently delivers a favorable trade-off across a range of $k$ values on MATH, AIME 2025 and AMC23. On MATH, W-REINFORCE matches the best Pass@$1$ score $76.6$ with PPO and has the highest performance for all $k\leq 64$, while still preserves competitive performance at $k=256$. On AIME 2025, W-REINFORCE performs even more strongly—achieving the best result across all $k$ values except $k=1$. These results confirm that W-REINFORCE, a simple weighted extension of the classic REINFORCE algorithm, strikes an effective balance between the strengths of PSR and NSR. By merely scaling down the weight of positive rewards, it achieves strong performance while preserving diversity. Despite its simplicity, W-REINFORCE consistently outperforms strong RL algorithms such as PPO and GRPO across most Pass@$k$, while the vanilla REINFORCE underperforms. These findings suggest that this simple variant of REINFORCE can serve as a competitive alternative to more complex RL algorithms when the model prior is strong (e.g., Qwen models). We provide an ablation study on the choice of $\lambda$ in Appendix˜E.
@@ 6 Related Work
Reinforcement learning has shown great promise in improving large language models, as demonstrated by the success of reinforcement learning from human feedback (RLHF) and from AI feedback (RLAIF), which align model responses with human preferences (lee2024rlaif, ; meng2024simpo, ; ouyang2022training, ; rafailov2023direct, ). More recently, reinforcement learning with verifiable rewards (dai2025s, ; gao2024designing, ; havrilla2024teaching, ; lambert2024t, ; liu2025understanding, ; shao2024deepseekmath, ; team2025intellect, ; wang2025reinforcement, ; xu2025scalable, ; yu2025dapoopensourcellmreinforcement, ; yuan2025vapo, ; yuan2025s, ; zhang2025srpo, ; wei2025truthrl, ) has attracted growing attention for its effectiveness in incentivizing reasoning in LLMs with rule-based rewards (fatemi2025concise, ; guo2025deepseek, ; kazemnejad2024vineppo, ; team2025kimi, ; zeng2025simplerl, ; dong2024rlhf, ; lambert2024rewardbench, ; mroueh@understandgrpo, ). Notably, DeepSeek-R1 (guo2025deepseek, ) and Kimi K1.5 (team2025kimi, ) demonstrate that RLVR can elicit emergent reasoning behaviors such as long chain of thought and self-reflection, and achieve strong performance across diverse reasoning tasks such as math and coding problems. Yeo et al. (yeo2025demystifying, ) further explore the emergence of long CoT across different RLVR setups.
Despite these advances, many prior works mainly focus on evaluating the model’s Pass@$1$ or greedy decoding performance, which might overlook the underlying change in model behavior (e.g., inference scaling performance). More importantly, the mechanisms behind RLVR for driving reasoning and generalization remain underexplored. In this work, we take a step further by decomposing the RLVR objective into two distinct learning paradigms—learning from correct responses and learning from incorrect responses. We analyze the learning signal at the token level through gradient analysis and evaluate the models extensively using Pass@$k$ with a wide spectrum of $k$. Our findings highlight the critical yet previously underappreciated role of negative rewards in RLVR.
@@
Inference-time scaling has emerged as a promising direction for enhancing model performance (balachandran2025inference, ; brown2024large, ; chen2024more, ; lightman2024lets, ; muennighoff2025s1, ; qu2025optimizing, ; setlur2025scaling, ; snell2024scaling, ; wang2024openr, ; wang2024math, ; wang2023selfconsistency, ; welleck2024decoding, ; wu2025inference, ; core23zhu, ; yao2023tree, ; zelikman2022star, ; zhao2025sample, ; wei2025webagent-r1, )—particularly in reasoning tasks where generating multiple candidate solutions (cobbe2021training, ; core23zhu, ) or longer reasoning traces (muennighoff2025s1, ; yeo2025demystifying, ; zeng2025revisiting, ) can help hit the correct answer. Through the lens of inference-time scaling, recent studies have raised questions about whether RL-trained models are better than the base models (yue2025does, ; hochlehnert2025sober, ). A recent work by (yue2025does, ) suggests that RLVR primarily adjusts the model’s output distribution toward high-reward responses,rather than eliciting new reasoning abilities beyond the base model. By adopting Pass@$k$ as metric, they find that RL-trained models have inferior inference-time scaling performance than the base models. These findings align with our findings that reinforcing correct samples hurts Pass@$k$ at larger $k$. Another concurrent work by (dang2025weight, ) studies supervised fine-tuning (SFT), and finds that diversity collapse during SFT adversely affects inference-time scaling performance. They show that SFT reduces generation diversity, leading to degraded Pass@$k$ performance, and that interpolating weights from early, potentially less overconfident checkpoints and later ones can effectively restore diversity and improve both Pass@$1$ and Pass@$k$.
@@ 7 Conclusion
In this work, we investigate the mechanism underlying RLVR for LM reasoning. By decomposing RLVR into positive and negative sample reinforcement, we reveal a surprising finding: solely penalizing incorrect samples can effectively enhance LM reasoning capabilities while preserving generation diversity. Experimental results show that NSR consistently improves performance across a wide Pass@$k$ spectrum and in many cases matches or outperforms strong RL algorithms such as PPO and GRPO. Our gradient analysis demonstrates that NSR works by suppressing incorrect responses and redistributing probability mass toward plausible alternatives based on the model prior. Building on these findings, we proposed a simple variant of REINFORCE, Weighted-REINFORCE, that upweights the negative sample reinforcement. Empirical results show that it achieves a good balance between PSR and NSR, and yields consistent Pass@$k$ improvements across multiple reasoning benchmarks. We discuss limitations and future work directions in Appendix˜F.
@@ Acknowledgments
We would like to thank the members of the Princeton NLP Group for their valuable feedback and discussions. We also thank the anonymous reviewers and the area chair for their time, effort, and constructive suggestions, which have greatly improved the quality of this paper. This research is partially funded by the National Science Foundation (IIS-2211779), the NVIDIA Academic Grant, the Lambda Research Grant, and the OpenAI Superalignment Fast Grant. MX is supported by the Apple Scholars in AI/ML PhD Fellowship.
@@ Appendix A Gradient Derivation
In summary, the gradient of Equation˜6 can be expressed as
$\frac{\partial\mathcal{L}}{\partial z_{v}}=\begin{cases}-R\cdot\pi_{v}\cdot(1-\pi_{v})&\text{if }v=y_{t}\quad\text{(sampled token)}\\ R\cdot\pi_{y_{t}}\cdot\pi_{v}&\text{if }v\neq y_{t}\quad\text{(unsampled token)}\end{cases}$
Appendix B NSR differs from Entropy and Unlikelihood Training in Gradients
@@
Gradient of unlikelihood training objective.
While NSR shares the goal of penalizing undesirable outputs with unlikelihood training [57, 26], their formulation and resulting dynamics are fundamentally different: NSR penalizes incorrect outputs using their raw probabilities, rather than minimizing $\log(1-\pi_{\theta}(\boldsymbol{y}|\boldsymbol{x}))$ as in unlikelihood training. This distinction results in different gradient behavior and training dynamics as shown below:
$\mathcal{L}_{\text{unlikelihood}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\log\big(1-\pi_{\theta}(y_{t}\mid\boldsymbol{x},\boldsymbol{y}_{<t})\big).$ (10)
Taking the derivative w.r.t. token logits $z_{v}$ yields:
[…]
Comparing Equation˜11 against Equation˜8, it can be seen that Equation˜11 induces a markedly different update: it pushes down the sampled token $y_{t}$’s logit and redistributes probability mass to other tokens in proportion to their current probabilities $\pi_{v}$. Consequently, confident predictions are penalized more aggressively, which may erode prior knowledge. In contrast, the gradient of NSR (Equation˜8) is dampened by a $(1-\pi_{v})$ factor. This has the opposite effect: as confidence in a sampled token grows, the gradient update shrinks when that token appears in an incorrect answer. By reducing the penalty on high-confidence tokens, NSR avoids destructive updates and effectively preserves pretrained knowledge.
Appendix C RLVR vs. RL with Reward Models
Different from RLVR, many RL setups require training reward models to assign scalar values that indicate the quality of generation (e.g., in RLHF [36]). During policy optimization, these rewards are often normalized within a batch by subtracting the mean, so that a sample’s final reward depends on its relative ranking compared to others in the same batch. As a result, a sample’s reward can change sign (from positive to negative or vice versa) depending on the overall quality of the batch, making it difficult to interpret rewards as absolute indicators of positive or negative samples. In contrast, RLVR applies a binary reward (+1+1/1-1). The signs of the rewards are inherently tied to the correctness of individual sequences, as each token in a sequence receives the same reward and the batch average reward always remains within [1,1][-1,1], thus reward normalization will preserve the original sign. This inherent sign preservation in RLVR is critical for our analysis, as it enables a clean and unambiguous separation into positive and negative learning paradigms based on the reward signs.
Appendix D Implementation Details
D.1 Objectives and Training Hyperparameters
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\beta\text{KL}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg),$
$\mathcal{L}_{\text{PSR}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\min\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},\text{clip}\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},1-\epsilon,1+\epsilon\right)\right),\qquad r(\boldsymbol{x},\boldsymbol{y})=1,$
$\mathcal{L}_{\text{NSR}}(\theta)=-\frac{1}{T}\sum_{t=1}^{T}\min\left(-\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},-\text{clip}\left(\frac{\pi_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})}{\pi_{\text{old}}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})},1-\epsilon,1+\epsilon\right)\right),\qquad r(\boldsymbol{x},\boldsymbol{y})=-1,$
@@
D.2 Prompt Templates
We adopt the prompt templates for Qwen models following [74], as shown in Tables˜2 and 3.
@@ D.2 Prompt Templates
D.3 Advantage Normalization
The implementation of computing advantage in verl includes an advantage normalization. However, for PSR and NSR, since we exclude the KL penalty, the advantage is equal to the raw reward (i.e., +1+1 for PSR and 1-1 for NSR). In this case, applying normalization would cause the advantage to be zero, thus eliminating the learning signal. As a result, we disable this normalization in PSR and NSR’s implementation. A recent work [61] also suggests that the reward normalization may not be necessary for certain settings.
Appendix E Impact of Reward Weighting Coefficient $\lambda$
To examine how sensitive W-REINFORCE is to the choice of $\lambda$, we conduct an ablation study with $\lambda\in{0,0.05,0.1,0.2,1}$ on the MATH dataset, with $\lambda=0$ corresponding to NSR and $\lambda=1$ corresponding to vanilla REINFORCE. As shown in Table˜4, performance is relatively stable when $\lambda\leq 0.2$, while larger values gradually degrade Pass@$k$ across all $k$. This confirms that the key to W-REINFORCE ’s effectiveness lies in moderately down-weighting the positive reward, which prevents the model from collapsing its distribution toward the highest-probability answers. When $\lambda=1$, Pass@$256$ drops substantially, underscoring that excessive emphasis on positive rewards harms exploration and diversity. Overall, these results suggest that W-REINFORCE is robust to the choice of $\lambda$, and a small $\lambda$ can offer a better trade-off between Pass@$1$ and Pass@$256$.
Table 4: Pass@$k$ results on MATH, AIME 2025 and AMC23 with different $\lambda$ using Qwen2.5-Math-7B. Bold and underlined numbers denote the best and second-best results for each $k$.
Pass@$k$
1 2 4 8 16 32 64 128 256
$\lambda$ MATH
0 75.7 82.4 86.9 90.1 92.4 94.1 95.3 96.2 96.9
0.05 75.6 82.2 86.7 89.9 92.2 94.0 95.3 96.3 97.1
0.1 76.6 82.8 87.1 90.2 92.4 94.1 95.3 96.1 96.7
0.2 75.8 81.2 85.2 88.3 90.7 92.6 94.0 95.1 95.9
1 74.8 79.1 82.4 84.9 86.9 88.5 89.9 91.1 92.0
$\lambda$ AIME 2025
0 10.0 14.6 19.2 24.1 29.3 34.6 40.2 46.0 53.3
0.05 9.0 13.6 18.3 22.9 27.8 33.3 39.4 46.2 53.3
0.1 10.6 15.3 20.0 24.7 29.7 34.6 40.5 47.8 56.7
0.2 9.4 13.2 17.3 21.8 26.5 31.1 35.7 39.6 43.3
1 9.2 12.5 16.3 21.0 26.4 32.3 38.6 44.7 50.0
$\lambda$ AMC23
0 60.9 70.0 77.4 83.2 87.6 91.1 94.5 97.9 100.0
0.05 61.6 70.9 77.8 82.9 87.1 91.2 95.1 98.1 100.0
0.1 62.0 70.0 77.0 83.1 87.8 91.8 95.2 97.1 97.5
0.2 60.4 68.0 73.9 79.3 84.5 89.5 93.6 96.4 97.5
1 61.8 68.5 73.4 77.0 80.8 85.1 89.3 91.9 92.5
Appendix F Limitations
@@
    Performance degradation with extended NSR training. We observe that extensive training with NSR (e.g., over hundreds of gradient steps) leads to a noticeable decline in performance, whereas W-REINFORCE does not exhibit this issue. However, such instability is not unique to NSR—standard RL algorithms like GRPO also experience similar degradation or training collapse under extended updates, as studied in recent works [30, 19, 79]. This suggests that NSR’s implicit mechanism for preserving prior knowledge may be insufficient to ensure stable training over the long term, and incorporating some degree of PSR may be necessary. Future work could explore more sophisticated methods for integrating NSR and PSR to design more robust RL algorithms, and investigate NSR as a potential warm-up phase before standard RL training given that it improves the full Pass@$k$ spectrum.